# MOOSE-Chem3: Toward Experiment-Guided Hypothesis Ranking via Simulated Experimental Feedback

### v2

@@ 3.1 Problem Formulation
that maps each hypothesis $h\in\mathcal{H}$ to a normalized performance score $s\in[0,1]$. The normalization provides a unified performance metric across heterogeneous research hypotheses and varying problem settings $q$, and can be defined relative to a domain-specific state-of-the-art benchmark established by experts.
@@ 3.2 Methodology
Figure 4: Experiment-guided ranking method.
We propose an experiment-guided ranking framework leveraging LLM agents, as illustrated in Figure 4. This design is informed by extensive consultations with chemistry domain experts, capturing key insights into hypothesis effectiveness.
These discussions identified a key criterion for hypothesis efficacy: effective hypotheses typically contain a sufficient number of key chemical components that collectively fulfill complementary mechanistic roles relevant to the research question $q$. Building upon this insight, our framework employs a structured, iterative approach comprising several distinct stages.
@@
Each candidate hypothesis $h\in\mathcal{H}$ is decomposed into its functional chemical components—distinct substructures or motifs potentially contributing to the target reaction mechanism. These components are then classified into three categories: effective, uncertain, and ineffective. Components deemed ineffective are excluded from further consideration to reduce computational overhead, as the initial hypothesis set $\mathcal{H}$ may yield a large number of components. The remaining components are clustered based on functional similarity, with each cluster representing a distinct mechanistic contribution to solving $q$. Individual elements within a cluster correspond to specific functional components, each traceable to its originating hypothesis $h$.
@@
Guided by the LLM’s prior chemical knowledge, the framework identifies the cluster most likely to contain components highly relevant to $q$. Within this selected cluster, the LLM agent further selects a hypothesis $h$ deemed most promising based on component relevance and prior understanding.
@@
The selected hypothesis $h$ is evaluated using the experimental executor (or simulator) $E$, yielding a normalized performance score $s$. The outcome of this experiment is then analyzed to evaluate the effectiveness of the chosen cluster and to validate or update the mechanistic assumptions made.
@@
Following each experimental evaluation, a detailed analysis is conducted, and the insights gained are integrated into a cumulative summary. This continually updated summary synthesizes insights from all prior analyses, highlighting effective clusters and guiding future hypothesis and cluster selections.
By iteratively leveraging prior chemical knowledge and empirical feedback, this framework systematically refines hypothesis prioritization. The overall objective is to efficiently identify the optimal hypothesis while minimizing the total number of experimental trials.
@@ 4.1 Simulator: Evaluating the Simulator with Real Experiment Results
To rigorously evaluate the performance of our simulator on advanced chemical problems, we curated a benchmark of 30 cutting-edge research questions, each associated with 3–6 mutually related candidate hypotheses, totaling 124 hypotheses. Ground-truth experimental outcomes were sourced from published literature, covering major subfields of chemistry, as detailed in § C.1
For each hypothesis, simulated results were generated using the proposed CSX-Sim and compared against the annotated experimental outcomes. This trend comparison is illustrated in § C.2. Evaluation focused on two key criteria: (1) trend alignment, measured by Spearman rank correlation, which assesses whether the predicted performances correctly reflect the relative ranking of ground-truth annotations within each research question. This criterion is critical, as hypothesis ranking primarily depends on relative performance differences; absolute offsets (e.g., uniform biases of ±0.2) have limited impact on ranking outcomes. Here we use “Perfect Consistency Indicator” (PCI) as metric, which quantifies the number of research questions for which the simulator achieves perfect trend alignment with experimental outcomes (2) predictive accuracy, quantified by root mean square error (RMSE), measuring absolute deviations between predicted and annotated performances. Detailed explanations of this metric and other predictive accuracy indicators are available in the § D. The comparative results are summarized in Table 1.
@@
We adopt the “Matched Score” (Yang et al., 2024b) as our primary baseline, which evaluates hypotheses by measuring their similarity to ground-truth references through a reference-based comparison. Additionally, we conduct two ablation studies on CSX-Sim to assess the contribution of its key components: (1) The first ablation (w/o CriticalPoints) disables the labeling of critical components $\mathcal{C}$, as defined in Equation 3, allowing hypotheses that lack essential components to still receive positive feedback from the simulator; (2) The second ablation (w/o ComponentExtraction) skips the extraction and weighting of critical components, directly computing mechanism similarity using prompts analogous to the final module in Figure 3.
@@
As shown in Table 1, CSX-Sim achieves superior performance across all metrics, with a Spearman correlation of 0.960, perfect consistency in 26 out of 30 questions, and the lowest RMSE of 0.213. Compared to the Matched Score baseline, CSX-Sim demonstrates substantial improvements in both trend alignment (+0.117 in Spearman) and robustness (+14 in PCI), while also reducing predictive error. Ablation studies further highlight the importance of critical component identification: removing CriticalPoints slightly degrades performance (Spearman 0.950, PCI 23/30), whereas omitting component extraction leads to significant drops in both alignment (Spearman 0.864) and accuracy (RMSE 0.272). These results underscore the necessity of fine-grained component analysis in achieving high-fidelity simulation feedback.
@@ 4.2 Experiment-Guided Ranking: Baselines and Ablation Study
We evaluate experiment-guided ranking on the TOMATO-chem dataset (Yang et al., 2024b), which includes 51 chemical problems, each annotated with a ground-truth (gdth) hypothesis. For each problem, we use the MOOSE-Chem framework (Yang et al., 2024b) to generate 63 additional candidate hypotheses that are distinct from the ground truth, resulting in 64 hypotheses per research question (1 gdth and 63 negatives).
To measure performance, we define the metric $N_{\text{trials}}$, representing the number of simulation-based evaluations required to identify the ground-truth hypothesis for each of the 51 problems. Lower values of $N_{\text{trials}}$ indicate more efficient hypothesis prioritization. Results are summarized in Table 2.
Method $N_{\text{trials}}$ (($\downarrow$( ↓)
Random Sampling 32.000
Pre-Experiment Ranking 33.280
@@
We consider two baselines: Pre-Experiment Ranking and Random Sampling. Pre-Experiment Ranking follows the strategy used in MOOSE-Chem (Yang et al., 2024b), where hypotheses are scored based on the model’s prior knowledge and ranked accordingly, without incorporating any experimental feedback. Random Sampling selects hypotheses uniformly at random, serving as a simple yet unbiased baseline.
As shown in Table 2, both baselines require over 32 trials on average to identify the ground-truth hypothesis, with Pre-Experiment Ranking (33.28 trials) slightly underperforming Random Sampling (32.00 trials). This counterintuitive result indicates that relying solely on prior model knowledge—without feedback—can lead to suboptimal prioritization, as initial estimation errors may mislead the ranking more than random choice.
@@ 4.3 Simulator: Ablation on Different $\phi(\cdot)$ with Different Levels of Distortion
To assess how simulator quality affects ranking performance, we leverage the observation that experiment-guided ranking is fundamentally an optimization process: navigating the hypothesis space to identify candidates with superior experimental performance. A high-fidelity simulator facilitates this search by providing informative feedback, while a degraded simulator misleads the process, making it harder to reach the optimum. Based on this perspective, we systematically introduce controlled distortions that worsen simulator fidelity from an optimization standpoint.
Specifically, we collaborated with chemistry PhD students to design three types of distortions commonly encountered in chemical research: local maxima/minima, plateaus, and cliffs. These noise patterns capture typical challenges in hypothesis evaluation, informed by domain expertise and heuristics. We defined three distortion levels—Simple Noise, Moderate Noise, and Complex Noise—and incorporated them into the hypothesis embedding function $\phi(\cdot)$ to simulate increasingly challenging feedback conditions. The composition and classification of constructed noise are detailed in § E
We evaluated CSX-Rank, CSX-Rank w/o Clustering, and CSX-Rank w/o Clustering & Analysis across three noise scenarios of increasing complexity: Simple Noise (3 maxima, 3 minima, 1 cliff, 2 plateaus), Moderate Noise (8 maxima, 12 minima, 3 cliffs, 4 plateaus), and Complex Noise (38 maxima, 22 minima, 4 cliffs, 4 plateaus).
As shown in Table 3, increasing noise complexity progressively degraded performance across all methods, as reflected by higher $N_{\text{trials}}$. CSX-Rank consistently outperformed its ablated variants, maintaining a substantial efficiency margin even under Complex Noise (32.7 vs. 36.5 and 40.5 trials). These results highlight the robustness of functional clustering and feedback analysis in mitigating misleading signals and preserving search efficiency. The findings align with Section 4.2, underscoring the critical role of each component in navigating noisy hypothesis spaces.
@@ 5 Related Work
Most prior work on hypothesis ranking has focused on pre-experiment ranking. Some approaches assign a score to each hypothesis and rank them accordingly, providing a simple and efficient solution (Yang et al., 2024a, b). Others adopt a pairwise ranking strategy, evaluating hypothesis pairs one at a time (Si et al., 2024; Liu et al., 2025). However, these methods rely solely on the internal reasoning of LLMs and do not incorporate feedback from experimental outcomes.
To our knowledge, few existing works leverage experimental feedback in hypothesis-driven tasks. Notably, recent methods in mathematics (Romera-Paredes et al., 2024; Shojaee et al., 2024) and programming (Qiu et al., 2024) incorporate feedback loops by refining hypotheses based on verification outcomes. These approaches rely on domains where extremely efficient verifiers are available, allowing for rapid hypothesis testing and direct refinement rather than explicit ranking. In contrast, our work focuses on natural science domains, where real experiments are significantly more costly, making such exhaustive trial-and-error strategies impractical. This difference motivates the need for a more deliberate experiment-guided ranking process, where each experiment must inform the prioritization of future hypotheses due to limited experimental bandwidth.
Roohani et al. (2024) address hypothesis generation in a genetic perturbation setting, where task-specific feedback can be directly computed (e.g., via gene overlap). In contrast, our work focuses on constructing general-purpose simulators for natural science domains, with an emphasis on chemistry due to the availability of annotated novel hypotheses from the literature (Yang et al., 2024b).
@@ 6 Conclusion
We present a systematic framework for experiment-guided hypothesis ranking in chemistry, addressing the critical challenge of limited access to real experimental feedback. By formalizing three foundational assumptions, we develop a high-fidelity simulator that approximates experimental outcomes based on hypothesis similarity, validated against a curated dataset of 124 hypotheses with reported wet-lab results. Building on this simulator, our proposed CSX-Rank method leverages functional clustering and iterative feedback analysis to efficiently prioritize hypotheses during the discovery process.
Empirical evaluations demonstrate that CSX-Rank significantly outperforms pre-experiment baselines, reducing the number of trials required to identify ground-truth hypotheses by more than 50% on the TOMATO-chem dataset. Ablation studies and controlled noise experiments further highlight the importance of analytical components and feedback integration for robust performance under increasingly challenging conditions.
Appendix A Extracting Key Chemical Components in the Simulator
A.1 A Framework for Extracting Critical Chemical Components in the Simulator
To better illustrate the specific framework of CSX-Sim for extracting key chemical components, as shown in Figure 5. For scientific hypotheses addressing specific problems, we categorize key chemical components and conclusions within the hypothesis. We then analyze the role and mechanism of each key chemical component based on the chemical problem and the conclusions drawn from the hypothesis. Finally, we review and output the key chemical components, their corresponding mechanisms, and the conclusions from the hypothesis.
@@ A.2 Prompt for Extracting Key Chemical Components in the Simulator
Each substance may be a key point. If it includes specific parameters like concentration or mass fraction (e.g., 0.3M NaCl, 10wtPVA), ensure these details are retained in the division process without losing specificity. If multiple substances are related and function together (e.g., potassium ferricyanide and potassium ferrocyanide as an oxidizing-reducing pair), group them as a single chemical key point based on their shared role or interdependence. Exclude elements from the scientific question that reappear in the hypothesis as prerequisites (e.g., if the question involves improving MXene nanosheets and the hypothesis enhances them with liquid metal, MXene nanosheets are a prerequisite, not a key point; liquid metal is the key point). Prerequisites should not be output or analyzed as key points. Distinguish key points from validation methods (e.g., elemental analysis to verify properties). Validation methods support the hypothesis but are not chemical key points. For each identified chemical key point, conduct a detailed and rigorous analysis of its role and function in relation to the scientific question. Use your chemical knowledge to explain the specific mechanism by which it addresses the problem, focusing on how it enhances the relevant properties or performance outlined in the question. Provide a clear, mechanistic explanation of its contribution and, if multiple key points exist, describe their interconnections.
@@
Example: Chemical Key Points 1. 10wtPVA (Polyvinyl Alcohol)

### v3

@@ 3.1 Task Formulation
that maps each hypothesis $h\in\mathcal{H}$ to a normalized performance score $s\in[0,1]$. The normalization provides a unified performance metric across heterogeneous hypotheses and varying problem settings $q$, and can be defined relative to a domain-specific state-of-the-art benchmark established by experts.
@@ 3.2 Methodology
Figure 4: Experiment-guided ranking policy within an in-context reinforcement learning framework.
Due to the high cost and data-scarce nature of wet-lab experiments in the natural sciences, conventional reinforcement learning (RL), which relies on extensive interaction and parameter updates, is often impractical. Our approach circumvents this bottleneck by formulating the learning process within the context window of a frozen large language model. This gradient-free, non-parametric paradigm relies solely on forward passes, enabling the agent to learn from minimal trials without costly fine-tuning. The framework leverages the LLM’s intrinsic reasoning capabilities, ensuring excellent generalizability to diverse scientific discovery tasks. Our agent, CSX-Rank, learns an optimal hypothesis selection policy via a formal sequential decision-making process.
We formulate experiment-guided ranking as a sequential decision-making process. At each timestep $t$, the agent observes a state $s_{t}$ representing the cumulative analysis of past experiments. It then performs an action $a_{t}$ by selecting a hypothesis $h\in H$ to test, receiving a reward $r_{t}$ from the experimental outcome. The trajectory is thus $\tau=[s_{0},a_{0},r_{0},\ldots]$.
Unlike standard RL settings that maximize cumulative reward, our objective is to identify the optimal hypothesis using the minimum number of experiments, reflecting the high cost of scientific exploration. The agent’s goal is to learn an optimal policy $\pi^{*}$ that minimizes the expected trials:
$\pi^{*}=\underset{\pi}{\operatorname*{arg\,min}}\,E_{\pi}[N_{\text{trials}}]$ (7)
Here, an effective policy $\pi(s_{t})$ leverages the accumulated knowledge in the state to make more strategic selections, thus minimizing $N_{\text{trials}}$. Our agent, CSX-Rank, implements this policy through the structured, iterative process detailed below (Figure 4).
@@
To generalize from specific results, the agent decomposes each hypothesis $h\in H$ into functional components, which are classified as effective, uncertain, or ineffective; the latter are pruned for efficiency. The remaining components are clustered by functional similarity, with each cluster representing a distinct mechanistic contribution to solving $q$. This yields a structured state representation $s_{t}$, where each element remains traceable to its originating hypothesis.
@@
To connect abstract mechanistic knowledge (clusters) with a concrete experiment (hypothesis), the policy $\pi(s_{t})$ selects the next action $a_{t}$ through a two-stage process. First, guided by prior domain knowledge, the LLM identifies the most promising cluster. Within this cluster, it then selects the most relevant hypothesis $h$, which defines the action $a_{t}$.
@@
The selected hypothesis $h$ (action $a_{t}$) is evaluated by the executor EE—either our high-fidelity simulator (CSX-Sim) or a real wet lab—which returns a normalized performance score $s\in[0,1]$. This score serves as the reward $r_{t}=E(a_{t})$, and its analysis quantifies the action’s success, grounding the policy in empirical results.
@@
To make learning cumulative, the agent integrates each experimental outcome into a running summary. This updated summary forms the new state $s_{t+1}$ for the next decision cycle, closing the RL loop and enabling the policy to refine systematically from prior knowledge and new feedback.
A key strength of our multi-step, component-driven framework is its inherent interpretability. By design, the agent must break down its decision process into an explicit, auditable trail—from extracting and clustering components to selecting the final hypothesis. Such structured and transparent reasoning is essential in scientific applications, allowing domain experts to examine the agent’s logic, build trust in its recommendations, and derive new insights.
@@ 4.1 Simulator: Evaluating the Simulator with Real Experiment Results
Table 1: Validating the simulator with collected experiment results from literature.
We curated a benchmark of 30 research questions and 124 hypotheses from published literature, each with experimentally validated outcomes spanning multiple domains (§ C.1). For each hypothesis, simulated results from CSX-Sim were compared against the annotated outcomes (§ C.2). Evaluation considered two criteria: (1) Trend alignment, measured by Spearman correlation, assessing whether predicted performances preserve the relative ordering of ground-truth outcomes. Because ranking depends on relative differences, we also report the Perfect Consistency Indicator (PCI), the number of questions with perfect alignment. (2) Predictive accuracy, measured by RMSE, capturing absolute deviations between predicted and experimental values (see § E for details and additional indicators). Comparative results are shown in Table 1.
@@
We adopt the “Matched Score” (Yang et al., 2025) as our primary baseline, which evaluates hypotheses by measuring their similarity to ground-truth references through a reference-based comparison. Additionally, we conduct two ablation studies on CSX-Sim to assess the contribution of its key components: (1) The first ablation (w/o CriticalPoints) disables the labeling of critical components $\mathcal{C}$, as defined in Equation 3, allowing hypotheses that lack essential components to still receive positive feedback from the simulator; (2) The second ablation (w/o ComponentExtraction) skips the extraction and weighting of critical components, directly computing mechanism similarity using prompts analogous to the final module in Figure 3.
@@
As shown in Table 1, CSX-Sim outperforms baselines across all metrics, demonstrating stronger trend alignment, greater robustness, and lower predictive error. Compared to the Matched Score baseline, it achieves notable gains in correlation and consistency while reducing error. Ablation studies confirm the importance of component analysis: removing CriticalPoints causes modest degradation, whereas omitting component extraction leads to substantial drops in alignment and accuracy. These results highlight the necessity of fine-grained component analysis for high-fidelity simulation feedback.
@@ 4.2 Experiment-Guided Ranking: Baselines and Ablation Study
We evaluate experiment-guided ranking on the TOMATO-chem dataset (Yang et al., 2025), which contains 51 scientific problems, each annotated with a ground-truth hypothesis. For each problem, the MOOSE-Chem framework (Yang et al., 2025) generates 63 additional candidates distinct from the ground truth, yielding 64 hypotheses per question (1 ground truth and 63 negatives). The disciplinary distribution is provided in § D. The dataset’s interdisciplinary nature, evident in its inclusion of topics from fields such as applied physics and biology, stems from its origin in scientific literature where "chemistry" papers are frequently co-labeled with other scientific fields. Performance is measured by $N_{\text{trials}}$, the number of simulation-based evaluations needed to identify the ground-truth hypothesis for each problem. Lower $N_{\text{trials}}$ indicates more efficient prioritization. Results appear in Table 2.
Method $N_{\text{trials}}$ $(\downarrow$)
Uninformed Search 32.500
Pre-Experiment Ranking 28.608
@@
We compare against two strategies: Uninformed Search and Pre-Experiment Ranking. Uninformed search selects hypotheses uniformly at random; Pre-experiment ranking scores hypotheses using only prior model knowledge, without feedback (Yang et al., 2025).
As shown in Table 2, Uninformed Search require over 32 trials on average, while Pre-Experiment Ranking reduces this to under 30—outperforming both naive baselines but still far behind CSX-Rank. This indicates that relying solely on prior knowledge yields only modest gains without feedback, whereas experiment-guided ranking substantially improves sample efficiency. A detailed scalability analysis is provided in § I.
@@ 4.3 Simulator: Ablation on Different $\phi(\cdot)$ with Different Levels of Distortion
To study how simulator fidelity affects ranking, we note that experiment-guided ranking is essentially an optimization process over hypothesis space. A high-fidelity simulator provides informative feedback to guide this search, while distortions mislead it. We therefore introduce controlled distortions into $\phi(\cdot)$ to simulate increasingly challenging feedback conditions.
In collaboration with domain experts, we designed three distortion types commonly observed in practice—local maxima/minima, plateaus, and cliffs—reflecting typical challenges in hypothesis evaluation. We further defined three distortion levels (Simple, Moderate, Complex), incorporating progressively more noise into $\phi(\cdot)$; full details appear in § F.
We evaluated CSX-Rank, CSX-Rank w/o Clustering, and CSX-Rank w/o Clustering & Analysis across these noise conditions. As shown in Table 3, higher noise complexity consistently degraded performance, increasing $N_{\text{trials}}$. Still, CSX-Rank outperformed its ablated variants, preserving a clear efficiency margin even under Complex Noise (32.7 vs. 36.5 and 40.5 trials). These results demonstrate the robustness of clustering and feedback analysis in mitigating misleading signals and maintaining search efficiency, aligning with Section 4.2.
@@ 5 Related Work
Most prior work on hypothesis ranking has focused on pre-experiment ranking. Some approaches assign a score to each hypothesis and rank them accordingly, providing a simple and efficient solution (Yang et al., 2024; 2025; Zhou et al., 2024). Others adopt a pairwise ranking strategy, evaluating hypothesis pairs one at a time (Si et al., 2024; Liu et al., 2025). However, these methods rely solely on the internal reasoning of LLMs and do not incorporate feedback from experimental outcomes.
To our knowledge, few existing works leverage experimental feedback in hypothesis-driven tasks, and those that do are confined to domains with highly efficient verifiers, enabling rapid hypothesis testing and direct refinement rather than explicit ranking. Notably, recent methods in mathematics (Romera-Paredes et al., 2024; Shojaee et al., 2024; Ma et al., 2024) and programming (Novikov et al., 2025; Qiu et al., 2024) incorporate feedback loops by refining hypotheses based on verification outcomes. In contrast, our work targets natural science domains, where real experiments are far more costly, rendering such exhaustive trial-and-error strategies impractical. This motivates the need for a more deliberate experiment-guided ranking process, designed to maximize the information gained from each costly experiment when prioritizing future hypotheses. Roohani et al. (2024) explore hypothesis generation in a genetic perturbation setting, where task-specific feedback can be computed directly (e.g., via gene overlap). This remains a niche domain where efficient verifiers are available. By contrast, our work focuses on constructing general-purpose simulators, enabling the study of experiment-guided ranking in settings where real experiments are costly and feedback is scarce.
@@ 6 Conclusion
We introduced the task of experiment-guided ranking and addressed its central bottleneck—the lack of scalable experimental feedback—by proposing a simulator grounded in three domain-informed conceptual foundations. Validated against 124 hypotheses, the simulator enables systematic evaluation of ranking policies. Building on this, we developed an in-context reinforcement learning framework with a clustering-based agentic policy that significantly outperforms pre-experiment baselines. Together, the simulator and policy provide a toolkit for advancing feedback-driven hypothesis discovery, with potential impact across the natural sciences.
Ethics statement
This work aims to accelerate beneficial scientific discovery, and the authors have read and adhered to the ICLR Code of Ethics. We acknowledge the potential for dual-use applications inherent in this research, as well as potential biases from our literature-based datasets and simulator conceptual foundations. We believe the benefits of a transparent and auditable framework for research outweigh these risks and are committed to its responsible application. No human subjects were involved in this study.
Reproducibility Statement
All code and data supporting this research are publicly available in an anonymous repository, a link to which is provided on the first page of this paper.
Appendix A Extracting Key FUNCTIONAL Components in the Simulator
A.1 A Framework for Extracting Critical Functional Components in the Simulator
To better illustrate the specific framework of CSX-Sim for extracting key functional components, as shown in Figure 5. For specific scientific problems, we use chemistry as an example to categorize the key functional components and conclusions within the hypothesis. We then analyze the role and mechanism of each key chemical component based on the chemical problem and the conclusions drawn from the hypothesis. Finally, we review and output the key chemical components, their corresponding mechanisms, and the conclusions from the hypothesis.
@@ A.2 Prompt for Extracting Key Chemical Components in the Simulator
Each substance may be a key point. If it includes specific parameters like concentration or mass fraction (e.g., 0.3M NaCl, 10wt% PVA), ensure these details are retained in the division process without losing specificity. If multiple substances are related and function together (e.g., potassium ferricyanide and potassium ferrocyanide as an oxidizing-reducing pair), group them as a single chemical key point based on their shared role or interdependence. Exclude elements from the scientific question that reappear in the hypothesis as prerequisites (e.g., if the question involves improving MXene nanosheets and the hypothesis enhances them with liquid metal, MXene nanosheets are a prerequisite, not a key point; liquid metal is the key point). Prerequisites should not be output or analyzed as key points. Distinguish key points from validation methods (e.g., elemental analysis to verify properties). Validation methods support the hypothesis but are not chemical key points. For each identified chemical key point, conduct a detailed and rigorous analysis of its role and function in relation to the scientific question. Use your chemical knowledge to explain the specific mechanism by which it addresses the problem, focusing on how it enhances the relevant properties or performance outlined in the question. Provide a clear, mechanistic explanation of its contribution and, if multiple key points exist, describe their interconnections.
@@
Example: Chemical Key Points 1. 10wt% PVA (Polyvinyl Alcohol)